OpenStudy (anonymous):

If 4^(2-X)=3, then 4^(2x-1)=?

OpenStudy (anonymous):

6.9

OpenStudy (anonymous):

Thanks but how did you get it? :)

OpenStudy (anonymous):

x=7.111111

OpenStudy (anonymous):

Hi but can you explain it? Thanks!

OpenStudy (anonymous):

Break up the power. \[4^{2-x} = 4^{2} * 4^{-x} = 3\] \[4^{-x} = 3/16\] Take the log base 4 of each side. -x = -1.206 x = 1.206 Checking, \[4^{2-1.206} = 3.006\] So the x value is valid. Plugging in x for the other equation, \[4^{2(1.206) -1} = 7.08 = 7.1\] If you use the EXACT value, you get 7.1111...

OpenStudy (anonymous):

64/9=7.1111

OpenStudy (anonymous):

you can also use logarithm

OpenStudy (anonymous):

Yes, you could take the log base 4 of each side right away, but I did not know if he/she was used to polynomials as a power, so I decided to break up the power.

OpenStudy (anonymous):

we can also try If 4^(2-X)=3, then 4^(2x-1)=? log4^(2-X)=log3 (2-X)log4=log3 (2-X)=log3/log4 2-(log3/log4)=x now use this to the other eq 4^(2x-1)=? 4^[(2(2-log3/log4)-1)]=? =64/9=7.11111 ans

OpenStudy (anonymous):

whews my pc hanged on me....lol

OpenStudy (anonymous):

anyway, the answer to both is correct she/he can choose whatever they doing at school............. lol